Skip to content

Changes for LAr calibration processing

  • Multiply OFCb by 1.5 for HEC channels which are in the 1.65 or 1.75 eta slice, to account for errors in the calculation of the DAC2MeV factors in these regions. This will eventually be removed in favour of a proper recalculation of these factors
  • Fixing isHECChannel for digital trigger (
  • Adding two new bad channel flags for super cells, so the cause of the masking is a bit clearer
  • don't attempt to switch to another gain when assigning autocorrelation for FCAL in low-mu settings
  • Updates to various job options and CA configs
Edited by Pavol Strizenec

Merge request reports